se voren 6. A man weighs 160 lb and is 5 feet 8 inches tall. What is his height and weight in cm and kg? 7. a. Using the defi i need question 6 the most
Answer #1

6)

Conversion to be used:

1 Kg = 2.205 lb

1 feet = 12 inch

1 inch = 2.54 cm

Given:

weight = 160 lb

= 160/2.205 Kg

= 72.6 Kg

height = 5 feet and 8 inch

= 5*12 inch + 8 inch

= 68 inch

= 68*2.54 cm

= 173 cm

Answer:

height = 173 cm

weight = 72.6 Kg